Please help me with this. Marion wants to dilute 2 gal of a 28% soap solution to make a 8% solution. How much pure water should she add to make the 8% solution?
Found 2 solutions by Fermat, checkley71:
Answer by Fermat(136)

2 gall of 28% soap solution
This comprises of 28% of 2 gall = 0.56 gall of pure soap + 1.44 gal of pure water.
Let Marion add n gall of pure water.
Total volume = n+2 gall
Concentratiuon of soap = 0.56 gall/(n+2) gall and this is to equal 8% = 0.08
0.56/(n+2) = 0.08
0.56 = 0.08(n+2)
0.56 = 0.08n + 0.16
0.40 = 0.08n
8n = 40
n = 5
Ammount of water to be added is,
n = 5 gallons
=============

Answer by checkley71(8403)

(2x.28)=(X+2).08 or .56=.08X+.16 or .40=.08X or x=.40/.08 or X=5 gal.